OpenStudy (anonymous):

Dan poured 5 cups of water into a rectangular container measuring 8 inches by 6 inches by 7 centimeters. [1 cup = 14.44 cubic inches; 1 inch = 2.54 cm] Part A: What is the maximum volume of water that the rectangular container can hold? Show your work. (3 points) Part B: How many cubic inches of water was poured into the container? (3 points) Part C: What was the height of the water in the container? Show your work. (4 points) Measurement is given as length by width by height.

OpenStudy (shinalcantara):

Part A: Volume of the rectangular container \[V = length \times width \times height\] since the units in the given is a mix of inches and cm, convert 7 cm into inches \[V = 8" \times 6" \times 7cm (\frac{ 1" }{ 2.54cm })\]

OpenStudy (shinalcantara):

Part B: 5 cups of water was poured and each cup contains 14.44cu. inches of water the amount of water poured then is \[V_{poured} = 5~cups(\frac{ 14.44~cu.inches }{ 1~ CUP})\]

OpenStudy (anonymous):

A)85344.4cubic inches B)72cubic inches C)1.5inches

OpenStudy (shinalcantara):

Part C: \[V = length \times width \times height\] \[V_{water}= 5(14.44)\] \[5(14.44) = 8 \times 6 \times h\] \[h = \frac{ 5 \times 14.44 }{ 8 \times 6 }\]

OpenStudy (anonymous):

A cylindrical water tank with 14m Diameter and 13m height. What is the capacity of tank in terms of Litters ?? @shinalcantara

OpenStudy (shinalcantara):

\[V = \pi r^2 h\] r = diameter/2 r = 14m/2 r = 7m ------------- \[V = \pi (7m)^2(13m)\]

OpenStudy (anonymous):

B)72cubic inches @sampathk

OpenStudy (shinalcantara):

What are your answer for Parts A, B and C @extremearianna ?

OpenStudy (shinalcantara):

*answers

OpenStudy (anonymous):

I sent you a message. :-) @shinalcantara

OpenStudy (shinalcantara):

\[2167.74cm^3[\frac{ 1dm^3 }{ (10cm)^3 }] = 2.16774 dm^3\]